Parth (parthkohli):

I suck at using mathematical tools in physics. Especially integration. What should I keep in mind when I mathematically model a physics problem? Just for instance, here's one problem: A cone is taken and a wire is uniformly tightened around its curved surface such that there are \(n\) turns per slant height. The half angle of the cone is \(\theta = 45^{\circ}\). The wire carries a current \(I\). Find the magnetic field at the centre of the base.

Parth (parthkohli):

@ganeshie8

Parth (parthkohli):

Here is the Biot-Savart Law:\[d\vec B = \frac{\mu _0 }{4\pi }\dfrac{i (\vec{d l }\times\vec r)}{r^3}\]

Parth (parthkohli):

@Michele_Laino

OpenStudy (michele_laino):

we have to use the superposition principle applied to magnetic field

OpenStudy (michele_laino):